OpenStudy (anonymous):

Hi can someone walk me thru the steps: (1/x^2-7x+10)=(x/x-5)+(1/x-2) I know that x^2-7x+10 can be factored to: (x-2)(x-5) What to do next

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{1}{x^2-7x+10}=\frac{x}{x-5}+\frac{1}{x-2} }\) add the fractions on the right side (by finding the common denominator).

OpenStudy (solomonzelman):

(note though, that you will have two restrictions for x \(\large\color{black}{ \displaystyle x\ne2 }\) \(\large\color{black}{ \displaystyle x\ne5 }\) and if your solution is for some reason x=2 or x=5, then it is extraneous)

OpenStudy (solomonzelman):

add the fractions on the right hand side for me please.

OpenStudy (anonymous):

What would be the common denominator how do i add them is the common denominator x^2-7x+10

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

I still don't under I got (1/x^2-7x+10)=(x^2-x-5)/(x^2-7x+10)

OpenStudy (anonymous):

Sorry about that I meant I still don't understand it

OpenStudy (anonymous):

I tried cross multiplying but I am stuck

OpenStudy (anonymous):

Confused please help

OpenStudy (phi):

is the problem to solve for x ?

OpenStudy (phi):

\[ \frac{1}{x^2-7x+10}=\frac{x}{x-5}+\frac{1}{x-2} \\ \frac{1}{(x-5)(x-2)}=\frac{x}{x-5}+\frac{1}{x-2} \] multiply both sides by (x-5)(x-2) like this: \[ (x-5)(x-2)\cdot \frac{1}{(x-5)(x-2)}=(x-5)(x-2)\frac{x}{x-5}+(x-5)(x-2)\frac{1}{x-2} \] now "cancel" where you have the same thing in the top and bottom \[ \cancel{(x-5)(x-2)}\cdot \frac{1}{\cancel{(x-5)(x-2)}}=\cancel{(x-5)}(x-2)\frac{x}{\cancel{x-5}}+(x-5)\cancel{(x-2)}\frac{1}{\cancel{x-2}} \] all of that mess becomes \[ 1 = (x-2)x + x-5 \] or \[ 1 = x^2 -2x +x -5 \] can you finish ?

OpenStudy (anonymous):

Yes thanks I think i can